# Decode Neural signal as Speech

### v1

@@ Abstract
Decoding language from brain dynamics is an important open direction in the realm of brain-computer interface (BCI), especially considering the rapid growth of large language models. Compared to invasive-based signals which require electrode implantation surgery, non-invasive neural signals (e.g. EEG, MEG) have attracted increasing attention considering their safety and generality. However, the exploration is not adequate in three aspects: 1) previous methods mainly focus on EEG but none of the previous works address this problem on MEG with better signal quality; 2) prior works have predominantly used “teacher-forcing” during generative decoding, which is impractical; 3) prior works are mostly “BART-based” not fully auto-regressive, which performs better in other sequence tasks. In this paper, we explore the brain-to-text translation of MEG signals in a speech-decoding formation. Here we are the first to investigate a cross-attention-based “whisper” model for generating text directly from MEG signals without teacher forcing. Our model achieves impressive BLEU-1 scores of 60.30 and 52.89 without pretraining & teacher-forcing on two major datasets (GWilliams and Schoffelen). This paper conducts a comprehensive review to understand how speech decoding formation performs on the neural decoding tasks, including pretraining initialization, training & evaluation set splitting, augmentation, and scaling law.
@@ 2 Related Works
BLEU-N (%) normal-$\uparrow$ ROUGE-1 (%)normal-$\uparrow$ WER (%) normal-$\uparrow$
@@ 4.4 Generated Samples
(1) GT
Arthur replaced the tray onto the lacquered cherry table that stood in the center of his ordered chamber
NeuSpeech
The rental car records pointed to lacquered cherry-schedung and stood in the center of his ordered chamber in motel from Acres comb in?
NeuSpeech w/ tf
Arthur replaced thethe cartray onto thethe lacquered cherry-table inthat stood in thethe center of his ordered chamber in
NeuSpeech w/ pt
Arthur replaced the tray onto the lacquered cherry table to match from Acres.
NeuSpeech w/ jt
The rental car records pointed to Acres and stay in the center of his ordered chamber
eeg-to-text (Wang & Ji, 2022)
He said he had no idea how long it would take him to get there
eeg-to-text (Wang & Ji, 2022) w/ tf
was the old with the tablekeyered table table and had in the middle of the room table
(2) GT
My patience was long gone and I was back in the car to warming up when Acres tapped on the window and told me he had found whatever he was looking for
NeuSpeech
My patience was long gone and I was back in the car. But when I heard that many of you were looking for whatever it was, but what about this?
NeuSpeech w/ tf
My patience was long gone and I was back in athe car.to warming up when Acres wastapped on thethe window and openedtold me he had found whatever he was looking for.
NeuSpeech w/ pt
My patience was long gone and I was back in the car.
NeuSpeech w/ jt
My patience was long gone and I was back in the car. The window was looking good when Acres is ongoing on her car and she now has…
eeg-to-text (Wang & Ji, 2022)
He said he had no idea how long it would take him to get back home
eeg-to-text (Wang & Ji, 2022) w/ tf
father was running gone I was ready to the dark drive up Iacia me the window said me to was a a he was looking for
(3) GT
Very well said Arthur then take your story to that and let him defend you when no civilized publisher will approach you
NeuSpeech
Very well said Arthur! Tell me your story, and I’ll defend it from you when no civilized publisher will approach you.
NeuSpeech w/ tf
Very well said Arthur!then thattake your story to andthat and let him defend you and no civilized publisher will approach you when
NeuSpeech w/ pt
Very well said Arthur then take your story to that and let him defend you when no civilized publisher will approach you very well
NeuSpeech w/ jt
Very well said! It was that and we can make sure that it and not let him defend you when no civilized publisher will approach you
eeg-to-text (Wang & Ji, 2022)
It was the first time I had seen him in a long time and it made me very nervous
eeg-to-text (Wang & Ji, 2022) w/ tf
little then he the seat and the place go it take it he one man would you

### v2

@@ Abstract
Decoding language from brain dynamics is an important open direction in the realm of brain-computer interface (BCI), especially considering the rapid growth of large language models. Compared to invasive-based signals which require electrode implantation surgery, non-invasive neural signals (e.g. EEG, MEG) have attracted increasing attention considering their safety and generality. However, the exploration is not adequate in three aspects: 1) previous methods mainly focus on EEG but none of the previous works address this problem on MEG with better signal quality; 2) prior works have predominantly used “teacher-forcing” during generative decoding, which is impractical; 3) prior works are mostly “BART-based” not fully auto-regressive, which performs better in other sequence tasks. In this paper, we explore the brain-to-text translation of MEG signals in a speech-decoding formation. Here we are the first to investigate a cross-attention-based “whisper” model for generating text directly from MEG signals without teacher forcing. Our model achieves impressive BLEU-1 scores of 60.30 and 52.89 without pretraining & teacher-forcing on two major datasets (GWilliams and Schoffelen). This paper conducts a comprehensive review to understand how speech decoding formation performs on the neural decoding tasks, including pretraining initialization, training & evaluation set splitting, augmentation, and scaling law. Code is available at https://github.com/NeuSpeech/NeuSpeech1/tree/main.
@@ 2 Related Works
BLEU-N (%) $\uparrow$ ROUGE-1 (%$\uparrow$ WER (%) $\uparrow$
@@ 4.4 Generated Samples
(1) GT Arthur replaced the tray onto the lacquered cherry table that stood in the center of his ordered chamber
NeuSpeech The rental car records pointed to lacquered cherry-schedung and stood in the center of his ordered chamber in motel from Acres comb in?
NeuSpeech w/ tf Arthur replaced thethe cartray onto thethe lacquered cherry-table inthat stood in thethe center of his ordered chamber in
NeuSpeech w/ pt Arthur replaced the tray onto the lacquered cherry table to match from Acres.
NeuSpeech w/ jt The rental car records pointed to Acres and stay in the center of his ordered chamber
eeg-to-text (Wang & Ji, 2022) He said he had no idea how long it would take him to get there
eeg-to-text (Wang & Ji, 2022) w/ tf was the old with the tablekeyered table table and had in the middle of the room table
(2) GT My patience was long gone and I was back in the car to warming up when Acres tapped on the window and told me he had found whatever he was looking for
NeuSpeech My patience was long gone and I was back in the car. But when I heard that many of you were looking for whatever it was, but what about this?
NeuSpeech w/ tf My patience was long gone and I was back in athe car.to warming up when Acres wastapped on thethe window and openedtold me he had found whatever he was looking for.
NeuSpeech w/ pt My patience was long gone and I was back in the car.
NeuSpeech w/ jt My patience was long gone and I was back in the car. The window was looking good when Acres is ongoing on her car and she now has…
eeg-to-text (Wang & Ji, 2022) He said he had no idea how long it would take him to get back home
eeg-to-text (Wang & Ji, 2022) w/ tf father was running gone I was ready to the dark drive up Iacia me the window said me to was a a he was looking for
(3) GT Very well said Arthur then take your story to that and let him defend you when no civilized publisher will approach you
NeuSpeech Very well said Arthur! Tell me your story, and I’ll defend it from you when no civilized publisher will approach you.
NeuSpeech w/ tf Very well said Arthur!then thattake your story to andthat and let him defend you and no civilized publisher will approach you when
NeuSpeech w/ pt Very well said Arthur then take your story to that and let him defend you when no civilized publisher will approach you very well
NeuSpeech w/ jt Very well said! It was that and we can make sure that it and not let him defend you when no civilized publisher will approach you
eeg-to-text (Wang & Ji, 2022) It was the first time I had seen him in a long time and it made me very nervous
eeg-to-text (Wang & Ji, 2022) w/ tf little then he the seat and the place go it take it he one man would you